OpenStudy (anonymous):

a company installs 5000 light bulbs each with an average life of 500 hours standard deviation of 100 hours and distribution approximated by a normal curve. find the percentage of bulbs that can be expected tovlast the period of time. less than 690 hours

OpenStudy (dumbcow):

first, find Z-score \[z = \frac{x-mean}{std dev} = \frac{690-500}{100} = 1.9\] then use standard normal table to find probability of being less than 1.9 deviations from mean